OpenStudy (goformit100):

Identify the correct slope and y-intercept of the equation 5x - 3y = -6 slope = 5 over 3; y-intercept = (0, 2) slope = -5; y-intercept = (0, -6) slope = negative 3 over 5; y-intercept = (0, 2) slope = 5 over 3; y-intercept = (0, -6)

OpenStudy (goformit100):

@Nurali

OpenStudy (anonymous):

If you rewrite that equation in terms of y, you'll get y = 5/3 x + 2 and the slope is the coefficient along x and y-intercept is the value you get when you put x = 0.

OpenStudy (goformit100):

What will be the correct option ?

OpenStudy (anonymous):

Well, the coefficient along x is 5/3 and when you put x=0 you'll get y=2 so the first one is correct.

OpenStudy (nurali):

5x - 3y = -6 3y=5x+6 y=5/3x+2 y=mx+c m=5/3 c=2 so x=0 y intercept (0,c) y intercept (0,2) slope = 5 over 3; y-intercept = (0, 2)